# Designing Phase Sensitive Probes of Monopole Superconducting Order

### v1

@@ A.1 Evaluating first order Josephson coupling
in which $eV=\mu_{L}-\mu_{R}$ is the voltage bias, and $\phi=\phi_{R}-\phi_{L}$ is the overall $\mathrm{U}(1)$ phase difference of the superconducting order parameters. Note that in the main text, we work at zero bias, $eV=0$. Above, the correlation function is given by $\Phi_{\mathrm{ret}}(t-t^{\prime},\phi)={-i\Theta(t-t^{\prime})}\langle[A(t),A(t^{\prime})]\rangle$, in which
@@
is the current operator and $T(\mathbf{k},\mathbf{k}^{\prime})$ is the tunneling matrix. It is assumed that the two states located at the left and right sides satisfy $\{c_{L,\mathbf{k}^{\prime},\sigma^{\prime}},c_{R,\mathbf{k},\sigma}\}=\{c_{L,\mathbf{k}^{\prime},\sigma^{\prime}},c^{\dagger}_{R,\mathbf{k},\sigma}\}=0$. Although the above correlation function is derived from first order perturbation theory, it corresponds to a second order process in $T_{\sigma,\sigma^{\prime}}$, which can be regarded as the tunneling of a Cooper pair.
@@
$\Phi(i\omega,\phi)=-\sum_{\mathbf{k},\mathbf{k}^{\prime}}\sum_{\sigma,\sigma^{\prime},\rho,\rho^{\prime}}T_{\sigma,\sigma^{\prime}}(\mathbf{k},\mathbf{k}^{\prime})T_{\rho,\rho^{\prime}}(-\mathbf{k},-\mathbf{k}^{\prime})\times\frac{1}{\beta}\sum_{ip_{n}}\mathcal{F}_{R,\sigma,\rho}^{\dagger}(\mathbf{k},ip_{n})\mathcal{F}_{L,\sigma^{\prime},\rho^{\prime}}(\mathbf{k}^{\prime},ip_{n}-i\omega).$ (46)
Above, $p_{n}=\pi(2n+1)/\beta$ with $n\in\mathbb{Z}$ are the fermion Matsubara frequencies, and $\mathcal{F}_{L,\sigma^{\prime},\rho^{\prime}}(\mathbf{k},ip_{n}-i\omega)$ and $\mathcal{F}^{\dagger}_{R,\sigma,\rho}(\mathbf{k},ip_{n})$ are the Fourier transforms of the anomalous Green’s functions, $\mathcal{F}_{L,\sigma^{\prime},\rho^{\prime}}(\mathbf{k},\tau)=\langle\mathcal{T}_{\tau}c_{L,-\mathbf{k},\rho^{\prime}}(\tau)c_{L,\mathbf{k},\sigma^{\prime}}(0)\rangle$ and $\mathcal{F}^{\dagger}_{R,\sigma,\rho}(\mathbf{k},\tau)=\langle\mathcal{T}_{\tau}c^{\dagger}_{R,\mathbf{k},\sigma}(\tau)c^{\dagger}_{R,-\mathbf{k},\rho}(0)\rangle$, with $\mathcal{T}_{\tau}$ being the time-ordering operator. It follows that the correlation function can be expressed as [31]
$\Phi(i\omega,\phi)=-\frac{1}{\beta}\sum_{ip_{n}}\sum_{\mathbf{k},\mathbf{k}^{\prime}}\mathrm{Tr}\left[\mathcal{F}_{L}(\mathbf{k},ip_{n}-i\omega)T^{\mathrm{T}}(-\mathbf{k},-\mathbf{k}^{\prime})\mathcal{F}^{*}_{R}(\mathbf{k},ip_{n})T(\mathbf{k},\mathbf{k}^{\prime})\right],$ (47)
in which the trace is taken over any internal degrees of freedom (spin, orbital, etc.).
@@
$\mathcal{H}_{\mathrm{BdG},\alpha}(\mathbf{k})=\left(\begin{array}[]{cc}\mathcal{H}_{\mathrm{kin},\alpha}(\mathbf{k})&\Delta_{\alpha}(\mathbf{k})e^{i\phi_{\alpha}}\\ \Delta^{\dagger}_{\alpha}(\mathbf{k})e^{-i\phi_{\alpha}}&-\mathcal{H}^{T}_{\mathrm{kin},\alpha}(-\mathbf{k})\end{array}\right),$ (48)
with $\alpha=L(R)$ corresponding to the left (right) side. Here, $\mathcal{H}_{\mathrm{kin},\alpha}(\mathbf{k})$ is the band Hamiltonian, $\Delta_{\alpha}(\mathbf{k})$ is the pairing matrix, and $\phi_{\alpha}$ is the overall $\mathrm{U}(1)$ superconducting phase, taken to be uniform. The anomalous Green’s functions can generally be expressed as [45]
@@
in which $\mathcal{G}_{\mathrm{kin},\alpha}(\mathbf{k},ip_{n})=[ip_{n}-\mathcal{H}_{\mathrm{kin}}(\mathbf{k})]^{-1}$ is the single-particle Green’s function of the band Hamiltonian in the absence of superconductivity. We work in the weak-coupling regime and consider one of the two following scenarios, under which the anomalous Green’s functions greatly simplify.
In the first scenario, the band Hamiltonian is spin-independent and thus the Green’s functions $\mathcal{G}_{\mathrm{kin},\alpha}$ are diagonal in spin space. For simplicity, we consider the case in which the Hamiltonian is in the spin-1/2 representation; though, generalization to systems with other internal degrees of freedom is straightforward. Consider the pairing matrix given by $\Delta_{\alpha}(\mathbf{k})=(d_{\alpha,0}(\mathbf{k})\sigma_{0}+\mathbf{d}_{\alpha}(\mathbf{k})\cdot\bm{\sigma})i\sigma_{y},$ in which $d_{\alpha,0}$ is the spin singlet pairing and $\mathbf{d}_{\alpha}$ is the $d$-vector corresponding to spin triplet pairing. In this scenario, the anomalous Green’s functions simplify as
@@
include corrections from nonunitary pairing orders. Above, $\xi_{\alpha,\mathbf{k}}=\xi_{\alpha,-\mathbf{k}}$ is the band energy dispersion at the left ($\alpha=L$) and right ($\alpha=R$) side of the junction. $d_{\alpha,0}$ and $\mathbf{d}_{\alpha}$ are both functions of $\mathbf{k}$, which we have dropped for convenience of notation. When the pairing is unitary ($\Delta_{\alpha}^{\dagger}\Delta_{\alpha}\propto\sigma_{0}$), the anomalous Green’s functions take a simpler form,
@@
in which $E_{\alpha,\mathbf{k}}=\sqrt{\xi_{\alpha,\mathbf{k}}^{2}+\mathrm{Tr}[\Delta_{\alpha}^{\dagger}(\mathbf{k})\Delta_{\alpha}(\mathbf{k})]/2}$ is the unperturbed dispersion of the BdG quasiparticle.
@@
(I) The first term, $e^{-i\phi}$, includes superconducting phase difference, $\phi=\phi_{R}-\phi_{L}$, and corresponds to the first order $2\pi$-periodic Josephson current. For higher-order contributions to the Josephson current, it is necessary to consider terms arising from multiple scatterings at the junction interface.
@@
in which $n_{F}$ is the Fermi-Dirac distribution. The final expression for the retarded correlation function is obtained via analytic continuation, $i\omega\rightarrow eV+i\delta$ as $\delta\rightarrow 0^{+}$.
@@
A.2 Effective spin-dependent tunneling in the presence of bulk spin-orbit interaction
We derive the tunneling matrix in the band representation, as introduced in Sec. II in the main text. The effective spin-orbit interaction arises from the tunneling of band eigenstates, $H_{\mathrm{link}}=\sum_{\mathbf{k},s;\mathbf{k}^{\prime},s^{\prime}}T_{s,s^{\prime}}^{(b)}(\mathbf{k},\mathbf{k}^{\prime})\psi^{\dagger}_{R,\mathbf{k},s}\psi_{L,\mathbf{k}^{\prime},s^{\prime}}+\mathrm{h.c.}$ in Eq. (11). Here, $\psi_{L(R),\mathbf{k},s}=\sum_{\sigma=\uparrow,\downarrow}[U_{L(R)}^{\dagger}(\mathbf{k})]_{s,\sigma}c_{L(R),\mathbf{k},\sigma}$ are the band eigenstates of the left (right) side in the absence of superconductivity, in which $U_{L(R)}(\mathbf{k})=u_{L(R),0}(\mathbf{k})\sigma_{0}+\mathbf{u}_{L(R)}(\mathbf{k})\cdot\bm{\sigma}$ is the unitary transformation to the band-diagonal representation. The tunneling matrix in the band-diagonal representation is expressed as $T^{(b)}(\mathbf{k},\mathbf{k}^{\prime})=T^{(b)}_{0}(\mathbf{k},\mathbf{k}^{\prime})+\mathbf{T}^{(b)}(\mathbf{k},\mathbf{k}^{\prime})\cdot\bm{\sigma}$, in which
@@ A.2 Effective spin-dependent tunneling in the presence of bulk spin-orbit interaction
Appendix B Effective pairing channels
We derive the effective monopole pairing order in the weak-coupling limit upon projecting to the helical band eigenstates and show the effective pairing channels inherited from the nontrivial spin texture of the Fermi surfaces. Consider the general Hamiltonian of a monopole superconductor, given in Eq. (13):
@@ Appendix B Effective pairing channels
In the weak-coupling regime, we project to the states at the Fermi surface, so that the band projected pairing matrix is given by
@@
Upon transforming back to the original representation, the projected effective pairing order reads
$\Delta^{(p)}_{\mathrm{inter}}(\mathbf{k})=U_{1}(\mathbf{k})\Delta^{(bp)}_{\mathrm{inter}}(\mathbf{k})U_{2}^{\mathrm{T}}(-\mathbf{k})\equiv{P}_{1}(\mathbf{k})\Delta_{\mathrm{inter}}(\mathbf{k}){P}_{2}^{\mathrm{T}}(-\mathbf{k}),$ (61)
in which ${P}_{i=1,2}(\mathbf{k})\equiv U_{i}(\mathbf{k})P_{+}^{(b)}U_{i}^{\dagger}(\mathbf{k})$. For example, consider a spin-$1/2$ basis and suppose that band eigenstates at the Fermi surface $\mathrm{FS}_{i=1,2}$ are given by $|\chi_{i,+}(\mathbf{k})\rangle=(u_{i}(\mathbf{k}),v_{i}(\mathbf{k}))^{\mathrm{T}}$. It follows that the projection operator in the spin-$1/2$ representation is
@@
When the Fermi surface has nontrivial spin textures, this can lead to superconducting pairing in new spin channels which differ from that of the original inter-Fermi surface pairing.
@@
$U_{1}(\mathbf{k})$ $=\sqrt{2\pi}\left(\begin{array}[]{cc}\mathcal{Y}_{+{\frac{1}{2}};\frac{1}{2},-\frac{1}{2}}(\bm{\Omega}_{\tilde{k}})&\mathcal{Y}_{{-\frac{1}{2}};\frac{1}{2},-\frac{1}{2}}(\bm{\Omega}_{\tilde{k}})\\ -\mathcal{Y}_{+{\frac{1}{2}};\frac{1}{2},+\frac{1}{2}}(\bm{\Omega}_{\tilde{k}})&-\mathcal{Y}_{{-\frac{1}{2}};\frac{1}{2},+\frac{1}{2}}(\bm{\Omega}_{\tilde{k}})\end{array}\right)$ (63c)
$U_{2}(-\mathbf{k})$ $=\sqrt{2\pi}\left(\begin{array}[]{cc}\mathcal{Y}_{+{\frac{1}{2}};\frac{1}{2},-\frac{1}{2}}(\bm{\Omega}_{\tilde{k}})&\mathcal{Y}_{{-\frac{1}{2}};\frac{1}{2},-\frac{1}{2}}(\bm{\Omega}_{\tilde{k}})\\ \mathcal{Y}_{+{\frac{1}{2}};\frac{1}{2},+\frac{1}{2}}(\bm{\Omega}_{\tilde{k}})&\mathcal{Y}_{{-\frac{1}{2}};\frac{1}{2},+\frac{1}{2}}(\bm{\Omega}_{\tilde{k}})\end{array}\right)$ (63f)
in which $\bm{\Omega}_{\tilde{\mathbf{k}}}$ is the spherical coordinate, with $\tilde{\mathbf{k}}=\mathbf{k}-\mathbf{K}_{0}$. The expressions for the half-integer monopole harmonics are given in Appendix C. When there is inter-Fermi surface $s$-wave pairing $\Delta_{\mathrm{inter}}(\mathbf{k})=\Delta_{0}i\sigma_{y}$, the projected pairing order in the spin-$1/2$ representation is given by
$\Delta^{(p)}_{\mathrm{inter}}(\mathbf{k})$
$=8\pi^{2}\Delta_{0}\left(\begin{array}[]{cc}-|\mathcal{Y}_{+\frac{1}{2};\frac{1}{2},-\frac{1}{2}}(\bm{\Omega}_{\tilde{k}})|^{2}\mathcal{Y}_{+\frac{1}{2};\frac{1}{2},-\frac{1}{2}}(\bm{\Omega}_{\tilde{k}})\mathcal{Y}_{-\frac{1}{2};\frac{1}{2},-\frac{1}{2}}(\bm{\Omega}_{\tilde{k}})&|\mathcal{Y}_{+\frac{1}{2};\frac{1}{2},-\frac{1}{2}}(\bm{\Omega}_{\tilde{k}})\mathcal{Y}_{+\frac{1}{2};\frac{1}{2},+\frac{1}{2}}(\bm{\Omega}_{\tilde{k}})|^{2}\\ -|\mathcal{Y}_{+\frac{1}{2};\frac{1}{2},-\frac{1}{2}}(\bm{\Omega}_{\tilde{k}})\mathcal{Y}_{+\frac{1}{2};\frac{1}{2},+\frac{1}{2}}(\bm{\Omega}_{\tilde{k}})|^{2}&-|\mathcal{Y}_{+\frac{1}{2};\frac{1}{2},+\frac{1}{2}}(\bm{\Omega}_{\tilde{k}})|^{2}\mathcal{Y}_{+\frac{1}{2};\frac{1}{2},+\frac{1}{2}}(\bm{\Omega}_{\tilde{k}})\mathcal{Y}_{-\frac{1}{2};\frac{1}{2},+\frac{1}{2}}(\bm{\Omega}_{\tilde{k}})\end{array}\right)$ (66)
@@
The effective pairing channels likewise can be seen in, for example, the tight-binding results discussed in Sec. IV.1.2 for a Josephson junction along the $z$ direction between an $s$-wave superconductor and $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}$ monopole superconductor. In the Josephson energy phase relation shown in Fig. 3 (b), states with momentum $k_{\parallel}=k_{\parallel,5}$ near $\tilde{k}_{F}$ have significant contributions to the first order Josephson tunneling, analogous to a junction between two $s$-wave superconductors. Similarly, the spin-singlet channel in the projected pairing order in Eq. (69) has maximal amplitude at the equator. In contrast, for states with momentum $k_{\parallel}=k_{\parallel,2}$ near the gap nodes, the first order Josephson current is suppressed while the second order $\pi$-periodic Josephson current is more significant. This can be attributed to the effective spin triplet channels in Eq. (69), which are more weighted relative to the spin singlet channel near the poles of the Weyl Fermi surface. Similar to a junction between a spin singlet and spin triplet superconductor with spin-independent tunneling at the junction link, this leads to higher order contributions to the Josephson current.

### v2

@@ A.1 First-order Josephson current and form factor in the presence of spin-dependent tunneling
where $eV=\mu_{L}-\mu_{R}$ is the voltage bias, and $\phi=\phi_{R}-\phi_{L}$ is the overall $\mathrm{U}(1)$ phase difference of the superconducting order parameters. Note that in the main text, we work at zero bias, $eV=0$. Here, the correlation function is given by $\Phi_{\mathrm{ret}}(t-t^{\prime},\phi)={-i\Theta(t-t^{\prime})}\langle[A(t),A(t^{\prime})]\rangle$, where
@@
is the current operator and $T(\mathbf{k},\mathbf{k}^{\prime})$ is the tunneling matrix. It is assumed that the two states located at the left and right sides satisfy $\{c_{L,\mathbf{k}^{\prime},\sigma^{\prime}},c_{R,\mathbf{k},\sigma}\}=\{c_{L,\mathbf{k}^{\prime},\sigma^{\prime}},c^{\dagger}_{R,\mathbf{k},\sigma}\}=0$. Although the above correlation function is derived from first-order perturbation theory, it corresponds to a second-order process in $T_{\sigma,\sigma^{\prime}}$, which can be regarded as the tunneling of a Cooper pair.
@@
$\Phi(i\omega,\phi)=$ $-\sum_{\mathbf{k},\mathbf{k}^{\prime}}\sum_{\sigma,\sigma^{\prime},\rho,\rho^{\prime}}T_{\sigma,\sigma^{\prime}}(\mathbf{k},\mathbf{k}^{\prime})T_{\rho,\rho^{\prime}}(-\mathbf{k},-\mathbf{k}^{\prime})$ (47)
×$\times\frac{1}{\beta}\sum_{ip_{n}}\mathcal{F}_{R,\sigma,\rho}^{\dagger}(\mathbf{k},ip_{n})\mathcal{F}_{L,\sigma^{\prime},\rho^{\prime}}(\mathbf{k}^{\prime},ip_{n}-i\omega).$
Here, $p_{n}=\pi(2n+1)/\beta$ with $n\in\mathbb{Z}$ are the fermion Matsubara frequencies, and $\mathcal{F}_{L,\sigma^{\prime},\rho^{\prime}}(\mathbf{k},ip_{n}-i\omega)$ and $\mathcal{F}^{\dagger}_{R,\sigma,\rho}(\mathbf{k},ip_{n})$ are the Fourier transforms of the anomalous Green’s functions, $\mathcal{F}_{L,\sigma^{\prime},\rho^{\prime}}(\mathbf{k},\tau)=\langle\mathcal{T}_{\tau}c_{L,-\mathbf{k},\rho^{\prime}}(\tau)c_{L,\mathbf{k},\sigma^{\prime}}(0)\rangle$ and $\mathcal{F}^{\dagger}_{R,\sigma,\rho}(\mathbf{k},\tau)=\langle\mathcal{T}_{\tau}c^{\dagger}_{R,\mathbf{k},\sigma}(\tau)c^{\dagger}_{R,-\mathbf{k},\rho}(0)\rangle$, with $\mathcal{T}_{\tau}$ being the time-ordering operator. It follows that the correlation function can be expressed as Sigrist and Ueda (1991)
$\Phi(i\omega,\phi)=$ $-\frac{1}{\beta}\sum_{ip_{n}}\sum_{\mathbf{k},\mathbf{k}^{\prime}}\mathrm{Tr}[\mathcal{F}_{L}(\mathbf{k},ip_{n}-i\omega)$ (48)
×$\times T^{\mathrm{T}}(-\mathbf{k},-\mathbf{k}^{\prime})\mathcal{F}^{*}_{R}(\mathbf{k},ip_{n})T(\mathbf{k},\mathbf{k}^{\prime})],$
where the trace is taken over any internal degrees of freedom (spin, orbital, etc.).
@@
$\mathcal{H}_{\mathrm{BdG},\alpha}(\mathbf{k})=\left(\begin{array}[]{cc}\mathcal{H}_{\mathrm{kin},\alpha}(\mathbf{k})&\Delta_{\alpha}(\mathbf{k})e^{i\phi_{\alpha}}\\ \Delta^{\dagger}_{\alpha}(\mathbf{k})e^{-i\phi_{\alpha}}&-\mathcal{H}^{\mathrm{T}}_{\mathrm{kin},\alpha}(-\mathbf{k})\end{array}\right),$ (49)
with $\alpha=L(R)$ corresponding to the left (right) side. Here, $\mathcal{H}_{\mathrm{kin},\alpha}(\mathbf{k})$ is the band Hamiltonian, $\Delta_{\alpha}(\mathbf{k})$ is the pairing matrix, and $\phi_{\alpha}$ is the overall $\mathrm{U}(1)$ superconducting phase, taken to be uniform. The anomalous Green’s functions can generally be expressed as Lüders and Usadel (1971)
@@
where $\mathcal{G}_{\mathrm{kin},\alpha}(\mathbf{k},ip_{n})=[ip_{n}-\mathcal{H}_{\mathrm{kin}}(\mathbf{k})]^{-1}$ is the single-particle Green’s function of the band Hamiltonian in the absence of superconductivity. We work in the weak-coupling regime and consider one of the two following scenarios, under which the anomalous Green’s functions greatly simplify.
In the first scenario, the band Hamiltonian is spin-independent and thus the Green’s functions $\mathcal{G}_{\mathrm{kin},\alpha}$ are diagonal in spin space. For simplicity, we consider the case where the Hamiltonian is in the spin-1/2 representation; though, generalization to systems with other internal degrees of freedom is straightforward. Consider the pairing matrix given by $\Delta_{\alpha}(\mathbf{k})=(d_{\alpha,0}(\mathbf{k})\sigma_{0}+\mathbf{d}_{\alpha}(\mathbf{k})\cdot\bm{\sigma})i\sigma_{y},$ where $d_{\alpha,0}$ is the spin singlet pairing and $\mathbf{d}_{\alpha}$ is the $d$-vector corresponding to spin triplet pairing. In this scenario, the anomalous Green’s functions are simplified as
@@
include corrections from nonunitary pairing orders. Here, $\xi_{\alpha,\mathbf{k}}=\xi_{\alpha,-\mathbf{k}}$ is the band energy dispersion at the left ($\alpha=L$) and right ($\alpha=R$) side of the junction. $d_{\alpha,0}$ and $\mathbf{d}_{\alpha}$ are both functions of $\mathbf{k}$, which we have dropped for convenience of notation. When the pairing is unitary ($\Delta_{\alpha}^{\dagger}\Delta_{\alpha}\propto\sigma_{0}$), the anomalous Green’s functions take a simpler form,
@@
where $E_{\alpha,\mathbf{k}}=\sqrt{\xi_{\alpha,\mathbf{k}}^{2}+\mathrm{Tr}[\Delta_{\alpha}^{\dagger}(\mathbf{k})\Delta_{\alpha}(\mathbf{k})]/2}$ is the unperturbed dispersion of the BdG quasiparticle.
@@
(I) The first term, $e^{-i\phi}$, includes superconducting phase difference, $\phi=\phi_{R}-\phi_{L}$, and corresponds to the first-order $2\pi$-periodic Josephson current. For higher-order contributions to the Josephson current, it is necessary to consider terms arising from multiple scatterings at the junction interface.
@@
where $n_{F}$ is the Fermi-Dirac distribution. The final expression for the retarded correlation function is obtained via analytic continuation, $i\omega\rightarrow eV+i\delta$ as $\delta\rightarrow 0^{+}$.
@@
A.2 Effective spin-dependent tunneling in the presence of bulk spin-orbit coupling
We derive the tunneling matrix in the band representation, as introduced in Sec. II in the main text. The effective spin-orbit interaction arises from the tunneling of band eigenstates, $H_{\mathrm{link}}=\sum_{\mathbf{k},s;\mathbf{k}^{\prime},s^{\prime}}T_{s,s^{\prime}}^{(b)}(\mathbf{k},\mathbf{k}^{\prime})\psi^{\dagger}_{R,\mathbf{k},s}\psi_{L,\mathbf{k}^{\prime},s^{\prime}}+\mathrm{h.c.}$ in Eq. (11). Here, $\psi_{L(R),\mathbf{k},s}=\sum_{\sigma=\uparrow,\downarrow}[U_{L(R)}^{\dagger}(\mathbf{k})]_{s,\sigma}c_{L(R),\mathbf{k},\sigma}$ are the band eigenstates of the left (right) side in the absence of superconductivity, where $U_{L(R)}(\mathbf{k})=u_{L(R),0}(\mathbf{k})\sigma_{0}+\mathbf{u}_{L(R)}(\mathbf{k})\cdot\bm{\sigma}$ is the unitary transformation to the band-diagonal representation. The tunneling matrix in the band-diagonal representation is expressed as $T^{(b)}(\mathbf{k},\mathbf{k}^{\prime})=T^{(b)}_{0}(\mathbf{k},\mathbf{k}^{\prime})+\mathbf{T}^{(b)}(\mathbf{k},\mathbf{k}^{\prime})\cdot\bm{\sigma}$, where
@@ A.2 Effective spin-dependent tunneling in the presence of bulk spin-orbit coupling
Appendix B Effective pairing in the presence of bulk spin-orbit coupling
We derive the effective monopole pairing order in the weak-coupling regime. Consider the general Hamiltonian of a monopole superconductor, given in Eq. (13):
@@ Appendix B Effective pairing in the presence of bulk spin-orbit coupling
In the weak-coupling regime, we project to the helical states at the Fermi surface, so that the band projected pairing matrix is given by
@@
Upon transforming back to the original representation in Eq. (59), the projected effective pairing order takes the following form:
$\Delta^{(p)}_{\mathrm{inter}}(\mathbf{k})$ $=U_{1}(\mathbf{k})\Delta^{(bp)}_{\mathrm{inter}}(\mathbf{k})U_{2}^{\mathrm{T}}(-\mathbf{k})$
$\equiv{P}_{1}(\mathbf{k})\Delta_{\mathrm{inter}}(\mathbf{k}){P}_{2}^{\mathrm{T}}(-\mathbf{k}),$ (62)
where ${P}_{i=1,2}(\mathbf{k})\equiv U_{i}(\mathbf{k})P_{+}^{(b)}U_{i}^{\dagger}(\mathbf{k})$. For example, consider a (pseudo)spin-$\uparrow$,$\downarrow$ , ↓ basis and suppose that band eigenstates at the Fermi surface $\mathrm{FS}_{i=1,2}$ are given by $|\chi_{i,+}(\mathbf{k})\rangle=(u_{i}(\mathbf{k}),v_{i}(\mathbf{k}))^{\mathrm{T}}$. It follows that the projection operator in the spin-$\uparrow$,$\downarrow$ , ↓ representation is
@@
$U_{1}(\mathbf{k})$ $=\sqrt{2\pi}\left(\begin{array}[]{cc}\mathcal{Y}_{+{\frac{1}{2}};\frac{1}{2},-\frac{1}{2}}(\bm{\Omega}_{\tilde{k}})&\mathcal{Y}_{{-\frac{1}{2}};\frac{1}{2},-\frac{1}{2}}(\bm{\Omega}_{\tilde{k}})\\ -\mathcal{Y}_{+{\frac{1}{2}};\frac{1}{2},+\frac{1}{2}}(\bm{\Omega}_{\tilde{k}})&-\mathcal{Y}_{{-\frac{1}{2}};\frac{1}{2},+\frac{1}{2}}(\bm{\Omega}_{\tilde{k}})\end{array}\right),$ (64c)
$U_{2}(-\mathbf{k})$ $=\sqrt{2\pi}\left(\begin{array}[]{cc}\mathcal{Y}_{+{\frac{1}{2}};\frac{1}{2},-\frac{1}{2}}(\bm{\Omega}_{\tilde{k}})&\mathcal{Y}_{{-\frac{1}{2}};\frac{1}{2},-\frac{1}{2}}(\bm{\Omega}_{\tilde{k}})\\ \mathcal{Y}_{+{\frac{1}{2}};\frac{1}{2},+\frac{1}{2}}(\bm{\Omega}_{\tilde{k}})&\mathcal{Y}_{{-\frac{1}{2}};\frac{1}{2},+\frac{1}{2}}(\bm{\Omega}_{\tilde{k}})\end{array}\right),$ (64f)
where $\bm{\Omega}_{\tilde{\mathbf{k}}}$ is the spherical coordinate, with $\tilde{\mathbf{k}}=\mathbf{k}-\mathbf{K}_{0}$. The expressions for the half-integer monopole harmonics are given in Appendix C. When there is inter-Fermi surface $s$-wave pairing $\Delta_{\mathrm{inter}}(\mathbf{k})=\Delta_{0}i\sigma_{y}$, the projected pairing order in the spin-$\uparrow$,$\downarrow$ , ↓ representation is given by
$\Delta^{(p)}_{\mathrm{inter}}(\mathbf{k})$ $=8\pi^{2}\Delta_{0}\left(\begin{array}[]{cc}-|\mathcal{Y}_{+\frac{1}{2};\frac{1}{2},-\frac{1}{2}}(\bm{\Omega}_{\tilde{k}})|^{2}\mathcal{Y}_{+\frac{1}{2};\frac{1}{2},-\frac{1}{2}}(\bm{\Omega}_{\tilde{k}})\mathcal{Y}_{-\frac{1}{2};\frac{1}{2},-\frac{1}{2}}(\bm{\Omega}_{\tilde{k}})&|\mathcal{Y}_{+\frac{1}{2};\frac{1}{2},-\frac{1}{2}}(\bm{\Omega}_{\tilde{k}})\mathcal{Y}_{+\frac{1}{2};\frac{1}{2},+\frac{1}{2}}(\bm{\Omega}_{\tilde{k}})|^{2}\\ -|\mathcal{Y}_{+\frac{1}{2};\frac{1}{2},-\frac{1}{2}}(\bm{\Omega}_{\tilde{k}})\mathcal{Y}_{+\frac{1}{2};\frac{1}{2},+\frac{1}{2}}(\bm{\Omega}_{\tilde{k}})|^{2}&-|\mathcal{Y}_{+\frac{1}{2};\frac{1}{2},+\frac{1}{2}}(\bm{\Omega}_{\tilde{k}})|^{2}\mathcal{Y}_{+\frac{1}{2};\frac{1}{2},+\frac{1}{2}}(\bm{\Omega}_{\tilde{k}})\mathcal{Y}_{-\frac{1}{2};\frac{1}{2},+\frac{1}{2}}(\bm{\Omega}_{\tilde{k}})\end{array}\right)$ (67)
@@
The effective pairing channels likewise can be seen in, for example, the tight-binding results discussed in Sec. IV.1.2 for a Josephson junction along the $z$ direction between an $s$-wave superconductor and $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}$ monopole superconductor. In the Josephson energy phase relation shown in Fig. 3(b), states with momentum $k_{\parallel}=k_{\parallel,5}$ near $\tilde{k}_{F}$ have significant contributions to the first-order Josephson tunneling, analogous to a junction between two $s$-wave superconductors. Similarly, the spin-singlet channel in the projected pairing order in Eq. (70) has maximal amplitude at the equator. In contrast, for states with momentum $k_{\parallel}=k_{\parallel,2}$ near the gap nodes, the first-order Josephson current is suppressed while the second-order $\pi$-periodic Josephson current is more significant. This can be attributed to the effective spin triplet channels in Eq. (70), which are more weighted relative to the spin singlet channel near the poles of the Weyl Fermi surface. Similar to a junction between a spin singlet and spin triplet superconductor with spin-independent tunneling at the junction link, this leads to higher-order contributions to the Josephson current.